# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks

### v1

@@ Abstract
Deep Neural Networks (DNNs) are increasingly deployed across a wide range of applications, from image classification to autonomous driving. However, ensuring their reliability remains a challenge and in many situations, alternative models with similar functionality and accuracy levels are available. Traditional accuracy-based evaluations often fail to capture behavioral differences between such models, particularly when testing datasets are limited, making it challenging to select or optimally combine models. Differential testing addresses this limitation by generating test inputs that expose discrepancies in the behavior of DNN models. However, existing differential testing approaches face significant limitations: many rely on access to model internals or are constrained by the availability of seed inputs, limiting their generalizability and effectiveness. In response to these challenges, we propose DiffGAN, a black-box test generation approach for differential testing of DNN models. Our method relies on a Generative Adversarial Network (GAN) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral discrepancies between models. Our method employs two custom fitness functions, one focused on diversity and the other on divergence, to guide the exploration of the GAN input space and identify discrepancies between the models’ outputs. By strategically searching the GAN input space, we show that DiffGAN can effectively generate inputs with specific features that trigger differences in behavior for the models under test.
@@
Deep Neural Network, Differential Testing, Test Generation, GAN, NSGA-II.
@@ 1 Introduction
In response to these challenges, we propose DiffGAN, a black-box test generation approach tailored for the differential testing of DNN models with similar accuracy levels on a given test dataset. By leveraging Generative Adversarial Networks (GANs) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II), DiffGAN generates diverse and valid triggering inputs that effectively reveal behavioral disagreements between the models under test. Since DiffGAN operates as a black-box method, it does not require access to the internal structure of the models or assume structural similarities between them, making it suitable for testing any pair of DNN models.
@@ 2.4 Differences from Adversarial Inputs
Adversarial samples are inputs subtly modified to mislead a specific single model by incorporating nearly imperceptible changes that may lead to incorrect predictions. Conversely, triggering inputs are designed to expose discrepancies between two versions of a model—for example the original and its compressed counterpart—by demonstrating that they produce divergent outcomes when fed with the same input. This distinction is crucial for understanding how models behave after modification or for comparing the behavior of two distinct models designed for the same task. Tian et al. [9] showed that adversarial samples of the original model rarely act as triggering inputs for compressed models. By using FGSM and CW adversarial methods [9] on the MNIST dataset, the authors found that only 18 out of 10,000 adversarial samples triggered the compressed models, supporting recent findings that model compression effectively defends against adversarial attacks. Also, several recent studies have shown that adversarial attacks are generally less effective at causing mispredictions in compressed models, as these models exhibit increased robustness against such input manipulations [15, 16]. We therefore exclude such kinds of inputs in our analysis.
@@ 3 Approach: Reformulation as an NSGA-II Search Problem
Since the models under test often exhibit similar high accuracy levels, GANs alone might not be effective in generating diverse triggering inputs that reveal behavioural disagreements between highly accurate models. To overcome these limitations, we leverage NSGA-II to effectively explore the large GAN’s latent space, thereby enhancing the generation of diverse triggering inputs. More specifically, our genetic search aims at finding latent vectors that maximize the generation of inputs produced by the GAN that are both diverse and reveal behavioral disagreements between pairs of DNN models under test. We also rely on NSGA-II since it is specifically adapted to our multi-objective search problem. It is widely used in the literature and has shown good performance in solving many search-based problems in DNN testing [25, 26, 27]. Many existing works in the literature have shown the effectiveness of NSGA-II in navigating the complex latent space of GANs and solving multi-objective optimization problems [28, 29, 30]. Unlike algorithms such as MOSA, the Many Objective Sorting Algorithm [31], which approaches each fitness function in isolation, NSGA-II tries to find solutions that offer diverse trade-offs among multiple objectives. More specifically, our search is driven by two objectives: (1) maximizing the difference between the output probability vectors of the inputs to generate triggering inputs, and (2) maximizing the diversity of generated inputs by the GAN model. We illustrate in Figure 1 our test generation approach and describe in the following the different steps of our approach along with the different components of our search algorithm.
Figure 1: Overview of DiffGAN
@@ 3.1 GAN Training Dataset
To effectively train our GAN model, we use the testing dataset available for the models under test. Indeed, access to the original training datasets of these models is often restricted in many real-world scenarios, particularly with proprietary models, hence our reliance on testing datasets.
To enrich the diversity of the GAN’s training data, we enhance the testing dataset of the models under test with five realistic image transformations, encompassing operations such as rotation, shifting, shearing, and adjustments to the images’ brightness and blurriness. These transformations are applied to every image in the dataset, following the same procedures and parameters established by Shorten et al. [32], which serve to ensure (1) the validity of the generated inputs and (2) labels preservation by these transformations. Consequently, the augmented dataset feeding into the GAN’s training process is expanded to five times its initial size, due to the application of five distinct image transformations on each original test input. Finally, we merge the original testing dataset with the augmented data to build the final GAN training dataset.
@@ 3.2 GAN Model Training
To effectively evaluate the performance of the generator, we used two SOTA metrics designed for evaluating GAN’s performance: the Fréchet Inception Distance (FID) and the Kernel Inception Distance (KID) [33]. The FID metric quantifies the difference between the distribution of GAN-generated images and the distribution of actual images in the training dataset. Note that in the literature [34], FID scores typically range from zero to a few hundred, with lower scores indicating higher quality in image generation, suggesting better GAN performance. The KID measures the similarity between the distributions of real and generated images by computing the maximum mean discrepancy between their feature representations [33]. Unlike FID, KID is calculated using a kernel function, making it more robust to variations in sample size. KID scores range from zero to one, the former reflecting superior quality in the images generated by the model.
@@
For each GAN model trained in our study, we generate a set of 10,000 images by inputting 10,000 randomly chosen latent vectors into the GAN. These generated images are then compared to a randomly selected subset of 10,000 real images from the GAN training dataset. To ensure the reliability of our results and reduce the impact of randomness, this procedure is conducted ten times for each GAN model. We then compute the average FID and KID scores across the ten iterations for each trained model. The trained GAN that obtains the lowest average scores for both FID and KID is selected as the optimal model.
@@ 3.3 Generation of Triggering Inputs
After training the GAN, we employ the NSGA-II search algorithm to effectively explore the GAN’s latent space, guiding the search towards the generation of diverse triggering input images. In this approach, a genetic algorithm manipulates and evolves a population of latent vectors through successive crossover and mutation operations. This evolutionary procedure iteratively refines these latent vectors towards the objectives. The selection process is guided by specific fitness functions, which, in our context, prioritize the selection of latent vectors that maximize the generation of diverse triggering input images. In other words, through successive iterations, we expect to increasingly produce test inputs that reveal a high number of diverse behavioral disagreements between the DNN models under test. To properly translate the generation process into a search problem using NSGA-II, we need to define the following elements.
@@ 3.3.1 Individuals and Initial Population
In genetic search, the initial population contains individuals consisting of a set of elements called genes. These genes are connected and form an individual also referred to as a solution. In our context, the initial population contains 100 latent vectors, each of size 100, as recommended in the literature [36]. We use random selection to build our initial population of individuals, thereby ensuring a diverse starting point for our genetic search.
@@ 3.3.2 Fitness Functions
This fitness function estimates the divergence between the output probability vectors of the models under test, when fed with the same input image generated by the trained GAN. More specifically, the fitness function takes as input a GAN latent vector $Z$ responsible for generating the candidate test input image. We feed the models under test with such image and compute the Euclidean distance between the resulting output probability vectors $Y_{m1}$ and $Y_{m2}$. The goal of maximizing this distance is to increase the probability of selecting latent vectors that lead to the generation of input images revealing significant behavioral disagreements between the models under test.
The Euclidean distance serves as a straightforward measure of the magnitude of discrepancy between model predictions, capturing how far apart the models’ outputs are in the probability space. The choice of Euclidean distance over other metrics, such as KL divergence [37], is motivated by its efficiency in terms of computation cost. However, the use of Euclidean distance or any other distance metrics alone has a limitation: they do not account for the differences in predicted labels across models. This means that while a high Euclidean distance indicates a discrepancy in the output probabilities, it does not necessarily reflect a disagreement in the models’ final predicted labels. This is particularly critical when the primary objective lies in searching for triggering inputs that lead to different label predictions by the models under test.
@@
Using a standard Euclidean distance in the context of differential testing is therefore not optimal and indicates the need to refine the fitness function, to ensure it not only considers the magnitude of input differences but also changes in the predicted labels, indicating different models predictions. To address this limitation, we modified the fitness function to reflect differences in model predictions. More specifically, given a GAN latent vector, if the models output different labels for a GAN-generated image, the fitness function increments the Euclidean distance by one to always assign a higher score for triggering input images. Otherwise, it returns the Euclidean distance between the two output probability vectors as illustrated in Equation 1:
@@
To illustrate the practical impact of our modified fitness function, reusing the examples above, Image 1, a triggering input, receives a fitness score of 1.3, whereas Image 2, a non-triggering input, is assigned a score of 0.3. This adjusted fitness function more accurately characterize inputs that are likely to reveal significant behavioral differences between models, prioritizing inputs that not only produce different probability vectors but also result in distinct classifications by the models.
@@
To reduce the computational complexity of comparing each generated image against all collected triggering inputs at each iteration of the search process, we approximate such distance by applying a clustering algorithm to group similar inputs. Then, we only return the minimum cosine distance between the generated image and the representative triggering inputs for all clusters. Preliminary experiments confirmed that, despite the cost of clustering, significant time could be saved in computing similarity new images, thus leading to generating more triggering inputs. More specifically, at each iteration of the search process, we apply the HDBSCAN [38] clustering algorithm on the set of the collected triggering input images to group similar inputs in one cluster. We select HDBSCAN for its ability to identify clusters of varying densities without requiring the prior specification of the number of clusters. This adaptability makes HDBSCAN particularly suited in our context.
Prior to clustering, we implement UMAP for dimensionality reduction, which assists in forming more distinct clusters [39, 27] at a lower cost. Subsequently, within each cluster, we identify the representative input—specifically, the input that has the maximum probability of association with its corresponding cluster222 https://hdbscan.readthedocs.io/en/latest/soft_clustering_explanation.html. The fitness function then calculates the diversity of the GAN-generated image by measuring the cosine distances between this image and the representatives of all clusters. Finally, the diversity-based fitness function returns the minimum of these cosine distances to quantify the diversity contribution of the image, as illustrated in Equation 2:
@@ 3.4.2 Mutation
Our genetic search also employs a mutation operator to further explore the search space. Following existing works that use genetic algorithms in GAN latent space [28, 40], we also adopt polynomial mutation [42] to address the continuous nature of the latent space. Polynomial mutation applies polynomial-distributed noise to the genes forming the selected individuals in our search. By applying such noise to the genes, it allows for the exploration of nearby solutions while still maintaining proximity to the original solution. This is also essential for efficiently navigating the complex and high-dimensional search space of GAN latent vectors. Polynomial mutation helps generate diverse solutions by introducing variability in the population of the search. This diversity is crucial for escaping local optima and finding globally optimal or near-optimal solutions.
@@ 4.1 Research Questions
RQ1. Do we generate more triggering inputs than the existing baseline approach with the same testing budget?
@@
Generating large numbers of triggering inputs is not sufficient. To the largest extent possible, they also need to be valid. We, therefore, aim to compare the validity of the triggering inputs produced by DiffGAN and those generated by DRfuzz. The concept of input validity is crucial for DNN testing in general and differential testing in particular since it ensures that the generated inputs are both relevant and meaningful. More specifically, they should be valid and belong to the targeted input domain. A valid input for a DNN model is defined as one that human experts within the domain can recognize and categorize accurately, using labels from the domain [3]. For example, in the context of handwritten digit recognition, an image is considered valid if it can be recognized as a digit by a human expert, despite potential imperfections. We, therefore, aim to compare in this research question the proportions of valid inputs that DiffGAN and DRfuzz generate. This comparison is crucial because generating a large number of triggering inputs is only beneficial if these inputs are to a large extent valid, as defined above.
@@
We aim in this research question to determine whether the triggering inputs generated by DiffGAN can more effectively guide the training of a model selection mechanism, allowing it to learn the conditions under which one DNN model outperforms another. In many real-world scenarios, it is common to encounter multiple DNNs that perform similarly in terms of overall accuracy, but exhibit significant variations in their performance under specific input conditions. This presents a challenge for developing a model selection mechanism that can dynamically choose the most suitable model based on the input characteristics without having to run all models in real-time. An effective solution to this problem is to use an ML-based selection mechanism, which can predict which of the available models will perform better for a given input. However, training such an ML-based model selection mechanism requires diverse and informative training data that captures and reveals a wide range of behavioural differences between the models. This is where the generation of triggering inputs becomes critical for the selection mechanism to more effectively learn under which condition one model is preferable over the other. Indeed, the effectiveness of the selection mechanism heavily depends on the quality of the triggering inputs used for training. Therefore, we aim to explore in this RQ whether DiffGAN can better guide the training process compared to DRfuzz. It is important to clarify that we do not intend to propose a novel ensemble method or voting mechanism in this work. Instead, this research question focuses on investigating an important application scenario, that is studying whether DiffGAN helps learn which model fares better under various input conditions and thus train a model selection mechanism.

### v2

@@ Abstract
Deep Neural Networks (DNNs) are increasingly deployed across a wide range of applications, from image classification to autonomous driving. However, ensuring their reliability remains a challenge and in many situations, alternative models with similar functionality and accuracy levels are available. Traditional accuracy-based evaluations often fail to capture behavioral differences between such models, particularly when testing datasets are limited, making it challenging to select or optimally combine models.
Differential testing addresses this limitation by generating test inputs that expose discrepancies in the behavior of DNN models. However, existing differential testing approaches face significant limitations: many rely on access to model internals or are constrained by the availability of seed inputs, limiting their generalizability and effectiveness. In response to these challenges, we propose DiffGAN, a black-box test generation approach for differential testing of DNN models. Our method relies on a Generative Adversarial Network (GAN) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral discrepancies between models. Our method employs two custom fitness functions, one focused on diversity and the other on divergence, to guide the exploration of the GAN input space and identify discrepancies between the models’ outputs. By strategically searching the GAN input space, we show that DiffGAN can effectively generate inputs with specific features that trigger differences in behavior for the models under test.
@@
Deep Neural Network, Differential Testing, Test Generation, Model Comparison, GAN, NSGA-II.
@@ 1 Introduction
In response to these challenges, we propose DiffGAN, a black-box test generation approach tailored for the differential testing of DNN models with similar accuracy levels. Our approach is specific to DNN models used in image classification, a domain in which such models are widely adopted. DiffGAN leverages Generative Adversarial Networks (GANs) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral disagreements between the models under test. The use of a GAN model for this purpose stems from its proven ability in generating realistic and novel images that conform to the original input domain. This is achieved by exploring the latent space of possible inputs. To effectively guide the generation process, we rely on NSGA-II to effectively search the GAN’s latent space and generate diverse triggering images. The genetic search employs two fitness functions, focusing on maximizing (1) diversity to avoid the generation of redundant triggering images, and (2) divergence to maximize the behavioral disagreements between the DNN models. Since DiffGAN operates as a black-box method, it does not require access to the internal structure of the models or assume structural similarities between them, making it suitable for testing any pair of DNN models. DiffGAN is focused on image classification problems, for which Deep Learning is widely applied across many applications.
@@ 2.4 Differences from Adversarial Inputs
Adversarial samples are inputs subtly modified to mislead a specific single model by incorporating nearly imperceptible changes that may lead to incorrect predictions. Conversely, triggering inputs are designed to expose discrepancies between two versions of a model—for example the original and its compressed counterpart—by demonstrating that they produce divergent outcomes when fed with the same input. This distinction is crucial for understanding how models behave after modification or for comparing the behavior of two distinct models designed for the same task. Tian et al. [9] demonstrated that, on the MNIST dataset, only 18 out of 10,000 adversarial samples generated using various adversarial attacks served as triggering inputs for differential testing. Similarly, prior studies [15, 16] indicate that adversarial inputs can be limited in their ability to expose differences between compressed models. Given that adversarial inputs target single model weaknesses and are not suitable for revealing differences in models’ behavior, they are excluded from our differential testing analysis.
@@ 3 Approach: Reformulation as an NSGA-II Search Problem
Since the models under test often exhibit similar high accuracy levels, GANs alone might not be effective in generating diverse triggering inputs that reveal behavioural disagreements between highly accurate models. To overcome these limitations, we leverage NSGA-II to effectively explore the large GAN’s latent space, thereby enhancing the generation of diverse triggering inputs. More specifically, our genetic search aims at finding latent vectors that maximize the generation of inputs produced by the GAN that are both diverse and reveal behavioral disagreements between pairs of DNN models under test.
We also rely on NSGA-II since it is specifically adapted to our multi-objective search problem. It is widely used in the literature and has shown good performance in solving many search-based problems in DNN testing [25, 26, 27]. Many existing works in the literature have shown the effectiveness of NSGA-II in navigating the complex latent space of GANs and solving multi-objective optimization problems [28, 29, 30]. Unlike algorithms such as MOSA, the Many Objective Sorting Algorithm [31], which approaches each fitness function in isolation, NSGA-II tries to find solutions that offer diverse trade-offs among multiple objectives. More specifically, our search is driven by two objectives: (1) maximizing the difference between the output probability vectors of the inputs to generate triggering inputs, and (2) maximizing the diversity of generated inputs by the GAN model. We illustrate in Figure 1 our test generation approach and describe in the following the different steps of our approach along with the different components of our search algorithm.
Figure 1: DiffGAN Framework Overview. The figure illustrates the two-phase process of DiffGAN. Phase 1 involves training a GAN using augmented test data to produce a realistic generator. After this, a transition occurs where the trained generator is carried forward into Phase 2, where NSGA-II is applied to search the generator’s latent space. Mutation and crossover operations, guided by divergence- and diversity-based fitness functions, iteratively evolve latent vectors to generate diverse and triggering inputs that expose behavioral discrepancies between DNN models.
@@ 3.1 GAN Training Dataset
To effectively train our GAN model, we use the testing dataset available for the models under test. Indeed, access to the original training datasets of these models is often restricted in many real-world scenarios, particularly with proprietary models, hence our reliance on testing datasets. To enrich the diversity of the GAN’s training data, we enhance the testing dataset of the models under test with five realistic image transformations, encompassing operations such as rotation, shifting, shearing, and adjustments to the images’ brightness and blurriness. These transformations are applied to every image in the dataset, following the same procedures and parameters established by Shorten et al. [32], which serve to ensure (1) the validity of the generated inputs and (2) labels preservation by these transformations. Consequently, the augmented dataset feeding into the GAN’s training process is expanded to five times its initial size, due to the application of five distinct image transformations on each original test input. Finally, we merge the original testing dataset with the augmented data to build the final GAN training dataset. We should note that training a separate GAN (or its generator) is not needed for each pair of models under test. The GAN is trained on the shared test dataset of the models under test, augmented with data transformations. Since the test dataset remains consistent across model pairs operating within the same domain, a single GAN can be reused for all model pairs tested within that domain.
@@ 3.2 GAN Model Training
To evaluate the generator’s performance effectively, we employed two SOTA metrics designed for evaluating GAN’s performance: the Fréchet Inception Distance (FID) and the Kernel Inception Distance (KID) [33]. The FID metric quantifies the difference between the distribution of GAN-generated images and the distribution of actual images in the training dataset. Note that in the literature [34], FID scores typically range from zero to a few hundred, with lower scores indicating higher quality in image generation, suggesting better GAN performance.
The KID measures the similarity between the distributions of real and generated images by computing the maximum mean discrepancy between their feature representations [33]. Unlike FID, KID is calculated using a kernel function, making it more robust to variations in sample size. KID scores range from zero to one, the former reflecting superior quality in the images generated by the model.
@@
During each iteration of the Optuna optimization process, a GAN model is trained, and 10,000 images are generated by feeding 10,000 randomly generated latent vectors into the trained GAN. These generated images are compared with a randomly selected subset of 10,000 real images from the GAN’s training dataset. To ensure the reliability of our results and reduce the impact of randomness, this process is repeated ten times per iteration. The average FID and KID scores across these ten runs are computed for the currently trained GAN model, and Optuna uses these scores to guide the search for the next set of hyperparameters, aiming to minimize FID and KID. Finally, the GAN model with the lowest average scores is selected as the optimal model.
@@ 3.3 Generation of Triggering Inputs
After training the GAN, we leverage its trained generator to produce triggering images. We employ the NSGA-II search algorithm to effectively explore the GAN’s latent space, guiding the search towards the generation of diverse triggering input images. In this approach, a genetic algorithm manipulates and evolves a population of latent vectors through successive crossover and mutation operations. This evolutionary procedure iteratively refines these latent vectors towards the objectives. The selection process is guided by specific fitness functions, which, in our context, prioritize the selection of latent vectors that maximize the generation of diverse triggering input images. In other words, through successive iterations, we expect to increasingly produce test inputs that reveal a high number of diverse behavioral disagreements between the DNN models under test. To properly translate the generation process into a search problem using NSGA-II, we need to define the following elements.
@@ 3.3.1 Individuals and Initial Population
In genetic search, the initial population contains individuals consisting of a set of elements called genes. These genes are connected and form an individual also referred to as a solution. In our context, the initial population contains 100 randomly generated latent vectors, each of size 100, as recommended in the literature [36]. Each dimension of a latent vector in the initial population is independently generated from a uniform distribution over the range $[-1,1]$, thereby ensuring a diverse and unbiased initialization in the search space.
@@ 3.3.2 Fitness Functions
This fitness function estimates the divergence between the output probability vectors of the models under test, when fed with the same input image generated by the trained GAN. More specifically, the fitness function takes as input a GAN latent vector $Z$ responsible for generating the candidate test input image. We feed the models under test with such image and compute the Euclidean distance between the resulting output probability vectors $Y_{m1}$ and $Y_{m2}$. The goal of maximizing this distance is to increase the probability of selecting latent vectors that lead to the generation of input images revealing significant behavioral disagreements between the models under test. The Euclidean distance serves as a straightforward measure of the magnitude of discrepancy between model predictions, capturing how far apart the models’ outputs are in the probability space. The choice of Euclidean distance over other metrics, such as KL divergence [37], is motivated by its efficiency in terms of computation cost.
However, the use of Euclidean distance or any other distance metrics alone has a limitation: they do not account for the differences in predicted labels across models. This means that while a high Euclidean distance indicates a discrepancy in the output probabilities, it does not necessarily reflect a disagreement in the models’ final predicted labels. This is particularly critical when the primary objective lies in searching for triggering inputs that lead to different label predictions by the models under test.
@@
Using a standard Euclidean distance in the context of differential testing is therefore not optimal and indicates the need to refine the fitness function, to ensure it not only considers the magnitude of input differences but also changes in the predicted labels, indicating different models predictions.
To address this limitation, we modified the fitness function to reflect differences in model predictions. More specifically, given a GAN latent vector, if the models output different labels for a GAN-generated image, the fitness function increments the Euclidean distance by one to always assign a higher score for triggering input images. Otherwise, it returns the Euclidean distance between the two output probability vectors as illustrated in Equation 1:
@@
To illustrate the practical impact of our modified fitness function, reusing the examples above, Image 1, a triggering input, receives a fitness score of 1.3, whereas Image 2, a non-triggering input, is assigned a score of 0.3. This adjusted fitness function more accurately characterizes inputs that are likely to reveal significant behavioral differences between models, prioritizing inputs that not only produce different probability vectors but also result in distinct classifications by the models.
@@
To reduce the computational complexity of comparing each generated image against all collected triggering inputs at each iteration of the search process, we approximate such distance by applying a clustering algorithm to group similar inputs. Then, we only return the minimum cosine distance between the generated image and the representative triggering inputs for all clusters. Preliminary experiments confirmed that, despite the cost of clustering, significant time could be saved in computing similarity scores of new images, thus leading to generating more triggering inputs. More specifically, at each iteration of the search process, we apply the HDBSCAN [38] clustering algorithm on the set of the collected triggering input images to group similar inputs in one cluster. We select HDBSCAN for its ability to identify clusters of varying densities without requiring the prior specification of the number of clusters. This adaptability makes HDBSCAN particularly suited in our context.
Prior to clustering, we implement UMAP for dimensionality reduction, which assists in forming more distinct clusters [39, 27] at a lower cost. Subsequently, within each cluster, we identify the representative input—specifically, the input that has the maximum probability of association with its corresponding cluster222 HDBSCAN Documentation—retrieved October 5, 2024, from https://hdbscan.readthedocs.io/en/latest/soft_clustering_explanation.html. The fitness function then calculates the diversity of the GAN-generated image by measuring the cosine distances between this image and the representatives of all clusters. Finally, the diversity-based fitness function returns the minimum of these cosine distances to quantify the diversity contribution of the image, as illustrated in Equation 2:
@@ 3.4.2 Mutation
Our genetic search also employs a mutation operator to further explore the search space. Following existing works that use genetic algorithms in GAN latent space [28, 40], we also adopt polynomial mutation [42] to address the continuous nature of the latent space.
Polynomial mutation applies polynomial-distributed noise to the genes forming the selected individuals in our search. By applying such noise to the genes, it allows for the exploration of nearby solutions while still maintaining proximity to the original solution. This is also essential for efficiently navigating the complex and high-dimensional search space of GAN latent vectors. Polynomial mutation helps generate diverse solutions by introducing variability in the population of the search. This diversity is crucial for escaping local optima and finding globally optimal or near-optimal solutions.
@@ 4.1 Research Questions
RQ1. Do we generate more triggering inputs than DRfuzz with the same testing budget?
@@
Generating large numbers of triggering inputs is not sufficient. To the largest extent possible, they also need to be valid. We, therefore, aim to compare the validity of the triggering inputs produced by DiffGAN and those generated by DRfuzz. The concept of input validity is crucial for DNN testing in general and differential testing in particular since it ensures that the generated inputs are both relevant and meaningful [3, 10]. More specifically, they should be valid and belong to the targeted input domain. A valid input for a DNN model is defined as one that human experts within the domain can recognize and categorize accurately, using labels from the domain [3]. For example, in the context of handwritten digit recognition, an image is considered valid if it can be recognized as a digit by a human expert, despite potential imperfections. We, therefore, aim to compare in this research question the proportions of valid inputs that DiffGAN and DRfuzz generate. This comparison is crucial because generating a large number of triggering inputs is only beneficial if these inputs are to a large extent valid, as defined above.
@@
We aim in this research question to determine whether the triggering inputs generated by DiffGAN can more effectively guide the training of a model selection mechanism, allowing it to learn the conditions under which one DNN model outperforms another. In many real-world scenarios, it is common to encounter multiple DNNs that perform similarly in terms of overall accuracy, but exhibit significant variations in their performance under specific input conditions [51, 52, 10, 11]. This presents a challenge for developing a model selection mechanism that can dynamically choose the most suitable model based on the input characteristics without having to run all models in real-time. An effective solution to this problem is to use an ML-based selection mechanism, which can predict which of the available models will perform better for a given input [51, 52]. However, training such an ML-based model selection mechanism requires diverse and informative training data that captures and reveals a wide range of behavioural differences between the models. This is where the generation of triggering inputs becomes critical for the selection mechanism to more effectively learn under which condition one model is preferable over the other. Indeed, the effectiveness of the selection mechanism heavily depends on the quality of the triggering inputs used for training. Therefore, we aim to explore in this RQ whether DiffGAN can better guide the training process compared to DRfuzz. It is important to clarify that we do not intend to propose a novel ensemble method or voting mechanism in this work. Instead, this research question focuses on investigating an important application scenario, that is studying whether DiffGAN helps learn which model fares better under various input conditions and thus train a model selection mechanism.